# The Effect of Massive Trans-Neptunian Objects in the Long-term Evolution and Leaking Rates of Neptune’s 3:2 and 2:1 Mean Motion Resonances

### v2

@@ 4.2 The evolution of the most massive Plutinos and Twotinos
Through our simulations, we identified which objects, characterized as the most massive resonant objects in section 2.3, were long-term resonant, meaning their resonant argument remained librating with a full amplitude below 360 over a 4 Gyr integration period. Here we should note two differences from the full data-set analysis: 1) For the massive sample, we can examine each object visually; therefore, we can distinguish very high amplitude librations from circulations. This distinction is difficult to implement automatically, so we limit the libration angle criterion to a fixed 340 in the analysis of the full data set. 2) After the first 10 Myr, the cadence of our data points is 400 times slower, and some very fast events could be missed.
@@
For the Twotinos, we found only two objects that are long-term resonant even in the massless DPs case, i.e., under the perturbations of the four GPs and the Sun. This highlights the importance of longer integrations for obtaining sounder classifications, for example, $2014\,{\rm DO}_{143}$, leaves the resonance very quickly, remains very close for 3.3 Gyr, and near the end of the simulation is recaptured into resonance; three more objects leave the resonance at $\sim$0.2, 2.8, and 2.8 Gyr, remaining as dropouts for the rest of the simulation; and the other four leave the resonance at $\sim$ 0.4, 0.6, 0.8, and 3.6 Gyr but these are shortly after ejected from the solar system.
@@
Finally, the addition of Pluto as a massive object in the integration, together with the ten massive Twotinos, destabilizes the set as compared to the simulation with massive Twotinos only, ejecting one more object after 4 Gyr. There is one additional object, $2005\,{\rm CA}_{79}$, that leaves the resonance after $3990$ Myr, and within 15 Myr (at $4005$ Myr is already on a hyperbolic orbit running away from the solar system). The effect is intermediate, with five objects in the resonance after 4 Gyr.
@@
4.3 Leaking Rates with and without Massive Plutinos and Twotinos
Resonances do not remain with constant populations throughout their history. Instead, a continuous leaking process keeps these populations in constant evolution, helping, for example, to feed the Jupiter Family comet population on secular time scales (e.g. W.-H. Ip & J. A. Fernandez, 1997; A. Morbidelli, 1997; M. A. Muñoz-Gutiérrez et al., 2019). Here, we analyze and compare the leaking rates of the resonances in the different scenarios considered. Simulations that include only the giant planets as gravitational perturbers serve as a point of comparison to simulations that include the ten most massive Plutinos and Twotinos, in addition to which we analyze the individual effect of Pluto on the evolution of both resonances, as well as together with the ten most massive Twotinos for the 2:1 MMR.
@@ 4.3 Leaking Rates with and without Massive Plutinos and Twotinos
The fractional evolution of each population for the different cases considered is shown in Figure 7. The two left panels of Fig. 7 correspond to the six scenarios studied for the 3:2 population, while the two right panels show the same six cases plus two scenarios of Pluto and the massive Twotinos for 2:1 populations. Furthermore, the upper two panels represent the SBDB dataset, while the bottom two represent the objects from the L7 model. Although the SBDB sample is relatively small, M. A. Muñoz-Gutiérrez et al. (2021) shows that, for orbits in chaotic phase space, each object evolves as a cloud of possibilities. By integrating the same orbit multiple times with different time steps, each particle unfolds into several possible evolutions. We use these results to enhance the statistical significance of the SBDB sample’s outcomes by running them three times (with time steps of 180, 200, and 220 days). The larger L7 samples make multiple runs more costly, while providing only minor insights. The resulting variations in the leaking rates due to the time-step variations are negligible for the 3:2 population, as indicated by the shaded areas surrounding the blue, red, and orange curves in the upper left panel. In contrast, for the 2:1 population, these errors appear to be more significant, causing some overlapping of the shaded areas (color-coded in the same way as for the 3:2 case) corresponding to the evolution of the SBDB population with and without massive Twotinos. We also present the cases, including Pluto alone with the 2:1 population, as well as Pluto together with the ten massive Twotinos (green curve, a case only applicable to the 2:1 populations). The lower panels show the same cases, color-coded in the same way as the upper panels, but for the L7 populations.
@@
Figure 7: Evolution of normalized counts of resonant objects from SBDB (upper panels) and L7 (lower panels) samples. The left panels correspond to simulations of the 3:2 population, and the right panels to the 2:1 population. The solid color lines indicate the number of objects, while the dotted lines correspond to exponential fits as described in equation 6 and table 3 in Section $4.$ Normalization is done with respect to the number of objects at 10 Myr. The colored shadows in the upper panels represent the variations obtained from the 3 different runs we performed in each scenario for the SBDB sample. Lines corresponding to the L7 model (bottom panels) do not have shadows, as their simulations were conducted only once due to computational expense.
In the case of Twotinos (right panels of Fig. 7), the evolutionary track of simulations that included only the giant planets is very similar for both the SBDB and L7 populations (blue curves in the upper and lower right panels). Adding the ten massive Twotinos into the simulations (orange curves) results in only a slight difference compared to the giant planet case alone, i.e., the perturbations produced by the then most massive Twotinos are almost negligible, as were the perturbations of the nine massive Plutinos besides Pluto for the 3:2 population. Surprisingly, a more critical instability on the 2:1 population is caused by Pluto alone, as its presence has a bigger effect than the ten massive Twotinos. This can be observed by comparing the red and green curves with the orange curve, and this effect is also evident in the L7 model, as shown by the corresponding curves in the bottom-right panel.
Pluto’s influence on the 2:1 populations may be attributed to the dwarf planet’s close encounters at aphelion with Twotino bodies. However, this suggests that Eris might also play an important role for Plutino bodies, given its perihelion is approximately 38 au. To explore this idea, we conducted numerical simulations using the 3:2 populations from the SBDB and L7 samples under perturbations from the giant planets and Eris. Our findings indicate that Eris has a minimal impact on the leak rate of the resonance, accounting for approximately 1% more leaking, when compared to the case of the giant planets alone. Another possibility for Pluto’s effect on the 2:1 population is their mutual 4:3 MMR. A more in-depth analysis of this phenomenon will be addressed in future work.
4.4 Leaking rate fittings and comparison with previous works
@@
where $N_{T}(t)$ is the resonant population at time $t$, which can be divided into two populations, $N_{S}$ and $N_{U}$ representing the stable and unstable populations, respectively, with $N_{U}^{0}$ being the unstable population at $t=0$. Note that $N_{S}$ is a constant offset related to the stable fraction of particles in the resonance, and $\tau$ gives the exponential decay rate of the unstable population. This behavior is consistent with the stochastic nature of the leaking phenomenon from resonances, i.e., the escape of objects from the resonances appears to be driven by random processes, as indicated by this distribution. At the same time, the constant offset can be attributed to hyper-stable objects, such as Pluto (e.g. T. Ito & R. Malhotra, 2025), but it consists of a significant fraction of the original population.
@@
Though not many, there have been some previous quantifications of the leaking rates from Neptune’s 3:2 and 2:1 MMRs. A. Morbidelli (1997) estimates a power law function of the time with an exponent of -0.5 for the 3:2 MMR, while M. S. Tiscareno & R. Malhotra (2009) found exponents of -0.55 and -0.77 for the 3:2 and 2:1 MMRs, respectively. Based on these works, S. Greenstreet et al. (2015) estimated the impact rates on Pluto and Charon during the last 4 Gyr, using a power law fitting given by:
@@
with $b=0.52$ for the 3:2 MMR and $b=0.77$ for the 2:1 MMR. S. Greenstreet et al. (2015) defined equation 8 so that the present time corresponds to 4.5 Gyr; therefore, to use it, we have to extrapolate it from 4.5 to 8.5 Gyr (i.e., we use it from the present time to 4 Gyr into the future).
In Fig. 8 we compare Eq. 8 with our exponential fittings for each resonance and for both the SBDB and L7 populations. Although S. Greenstreet et al. (2015) models only consider the four giant planets as perturbers, in this comparison, we include simulations with only the giant planets as perturbers, as well as our most complete simulations; that is, in the case of Plutinos, the case including the ten massive Plutinos, and for Twotinos, the case including the ten massive Twotinos and Pluto.
Overall, we only find agreement with our SBDB simulations without massive Plutinos and Twotinos (blue lines in both panels), having a very close match for the 2:1 MMR, and a good match to the initial slope in the 3:2 resonance. Since the real Trans-Neptunian region contains massive objects, it stands to reason that the erosion of the MMRs is faster than what our GP-only simulations or Greenstreet equations imply, suggesting that the MMRs were more heavily populated than the Greenstreet model predicts, at least for the past 1 Gyr (based on a short extrapolation of our equations into the past). This in turn has implications for the age of Pluto’s (and Charon’s) surface, with our model suggesting younger surfaces than predicted by S. Greenstreet et al. (2015).
Figure 8: Leaking rate fit comparisons. Our rates (solid colored lines) come from Table 3, but have been renormalized so they start with a value of 1.00 at $t=0$ Gyr. The dashed line corresponds to Eq. 8, taken from S. Greenstreet et al. (2015).
4.5 Effect of massive Plutinos and Twotinos on the evolution of resonant populations
@@ 5 Conclusions
In section 4.1, we conducted an analysis of the final state of the simulations that include only the GPs after 4 Gyr of integration. The sculpting of the orbital parameters tends to converge for the SBDB and the L7 populations, especially in semimajor axis. However, the median inclination of the 3:2 population, as well as the median eccentricity of the 2:1 population, did not really converge, indicating that no amount of evolution will make these distributions consistent among each other. The SBDB sample has observational biases, and whether there are inclination over or underestimations in the L7 model is beyond the scope of this paper. We also note that larger inclinations facilitate the escape from resonance, thus particles in the large-$i$ tail leak faster; this results in the median inclinations of all the populations being smaller at 4 Gyr.
The comparison between the initial and final distribution of the libration amplitudes and libration centers revealed that asymmetries in the leading and trailing populations of the 2:1 population remain unchanged over the evolution of our simulations: L7 objects, which start with a symmetrical distribution, end up with the same rough amount of leading and trailing objects, and the SBDB objects, which start with an uneven distribution among libration centers, ends up with a different, still uneven distribution at the 4 Gyr mark. This suggests that asymmetries in the leading and trailing populations probably arise due to a secondary mechanism at the early stage of the formation of the original populations, most commonly attributed to Neptune’s migration at the early stages of the Solar System (e.g., D. Nesvorný, 2018, and references therein).
By adding the massive perturbers to the simulations, as discussed in Section 4.2, we explored the effect of these additional perturbations among themselves, as well as over the resonant populations as a whole. Regarding the massive objects, from the ten most massive Plutinos, we found that when only the GPs are considered, all ten objects are resonant for the whole 4 Gyr, while adding mass results in 9 being resonant for 4 Gyr, while the remaining one, $2003\,{\rm UZ}_{413}$, is resonant for about 1.5 Gyr. For the massive Twotinos, even when only the GPs are included as perturbers, only two massive Twotinos are resonant for the whole 4 Gyr (although a third one is recaptured near the end of the simulation); the other 8 leave the resonance at different times between 20 Myr and 3.2 Gyr. When adding mass to these Twotinos, they become more stable, with six remaining resonant for the 4 Gyr (and one more being recaptured). Yet when adding Pluto as a massive TNO, the set becomes less stable with only four massive Twotinos remaining resonant for the 4 Gyr (and, once more, one more being recaptured).
@@
The addition of massive bodies into the simulations had a prominent effect on each group, drastically diminishing their stability in some cases, particularly for the 3:2 population. Within the 4 Gyr of our simulations, Pluto is responsible for more than 24% of the objects escaping the 3:2 population, while only 2% is attributed to the other nine most massive Plutinos. The ten most massive Twotinos create a difference of 4% in the number of objects leaving the resonance. The most surprising feature found is the influence of Pluto on the 2:1 population, where 13% of the escapes are caused by the addition of Pluto alone. The influence of Pluto over the stability of the 2:1 MMR is quite unexpected. It could be attributed to perturbations from the dwarf planet when it lies closer to aphelion, when its distance to objects trapped in 2:1 MMR could be drastically reduced, increasing the interaction probability, but if such were the case, Eris should be as relevant for the 3:2 population stability given its perihelion of $\sim 38.4$ au. Preliminary test simulations showed that Eris does not significantly perturb the 3:2 population; thus, a different possibility is that Pluto affects the 2:1 MMR populations through their mutual 4:3 MMR. A more in-depth study of this phenomenon is left for future work.
When comparing our results with those of previous authors, we found that simulations of the SBDB without massive perturbers are mostly consistent with extrapolations of the theoretical models (e.g. S. Greenstreet et al., 2015), however, our simulations including massive TNOs result in greater leaking rates.
@@
We acknowledge the referee, B. Gladman, for a careful reading and report that helped improve this paper. This research was performed using services/resources provided by Grid UNAM, which is a collaborative effort driven by DGTIC and the research institutes of Astronomy, Nuclear Science and Atmosphere Science and Climate Change at UNAM. M.A.M. acknowledges Universidad de Atacama for the DIUDA grant No. 88231R14. A.P.-V. acknowledges the DGAPA-PAPIIT grant IA103224. {contribution} MAMG and AP came up with the initial research concept, most of the writing, and most of the editing. SR was responsible for the computer simulations, as well as some writing and editing. APV and CP contributed to reviewing, writing, and editing.

### v3

@@ IV.2 The evolution of the most massive Plutinos and Twotinos
Through our simulations, we identified which objects, characterized as the most massive resonant objects in section II.3, were long-term resonant, meaning their resonant argument remained librating with a full amplitude below 360 over a 4 Gyr integration period. Here we should note two differences from the full data-set analysis: 1) For the massive sample, we can examine each object visually; therefore, we can distinguish very high amplitude librations from circulations. This distinction is difficult to implement automatically, so we limit the libration angle criterion to a fixed 340 in the analysis of the full data set. 2) After the first 10 Myr, the cadence of our data points is 400 times slower, and some very fast events could be missed.
@@
For the Twotinos, we found only two objects that are long-term resonant even in the massless DPs case, i.e., under the perturbations of the four GPs and the Sun. This highlights the importance of longer integrations for obtaining sounder classifications, for example, $2014\,{\rm DO}_{143}$, leaves the resonance very quickly, remains very close for 3.3 Gyr, and near the end of the simulation is recaptured into resonance; three more objects leave the resonance at $\sim$0.2, 2.8, and 2.8 Gyr, remaining as dropouts for the rest of the simulation; and the other four leave the resonance at $\sim$ 0.4, 0.6, 0.8, and 3.6 Gyr but these are shortly after ejected from the solar system. It is worth noting that additional observations would also be beneficial in constraining the orbital elements; this is especially true for the Twotinos, for which having accurate orbits may be critical due to the complex structure of their resonant phase-space.
@@
Finally, the addition of Pluto as a massive object in the integration, together with the ten massive Twotinos, destabilizes the set as compared to the simulation with massive Twotinos only, ejecting one more object after 4 Gyr. There is one additional object, $2005\,{\rm CA}_{79}$, that leaves the resonance after $3990$ Myr, and within 15 Myr (by $4005$ Myr, it is already on a hyperbolic orbit running away from the solar system). The effect is intermediate, with five objects in the resonance after 4 Gyr.
@@
IV.3 Leaking Rates with and without Massive Plutinos and Twotinos
Resonances do not remain with constant populations throughout their history. Instead, a continuous leaking process keeps these populations in constant evolution, helping, for example, to feed the Jupiter Family comet population on secular time scales (e.g. Ip and Fernandez, 1997; Morbidelli, 1997; Muñoz-Gutiérrez et al., 2019). Here, we analyze and compare the leaking rates of the resonances in the different scenarios considered. Simulations that include only the giant planets as gravitational perturbers serve as a point of comparison to simulations that include the ten most massive Plutinos and Twotinos, in addition to which we analyze the individual effect of Pluto on the evolution of both resonances, as well as together with the ten most massive Twotinos for the 2:1 MMR.
@@ IV.3 Leaking Rates with and without Massive Plutinos and Twotinos
The fractional evolution of each population for the different cases considered is shown in Figure 7. The two left panels of Fig. 7 correspond to the six scenarios studied for the 3:2 population, while the two right panels show the same six cases plus two scenarios of Pluto and the massive Twotinos for 2:1 populations. Furthermore, the upper two panels represent the SBDB dataset, while the bottom two represent the objects from the L7 model. Although the SBDB sample is relatively small, Muñoz-Gutiérrez et al. (2021) shows that, for orbits in chaotic phase space, each object evolves as a cloud of possibilities. By integrating the same orbit multiple times with different time steps, each particle unfolds into several possible evolutions. We use these results to enhance the statistical significance of the SBDB sample’s outcomes by running them three times (with time steps of 180, 200, and 220 days). The larger L7 samples make multiple runs more costly, while providing only minor insights. The resulting variations in the leaking rates due to the time-step variations are negligible for the 3:2 population, as indicated by the shaded areas surrounding the blue, red, and orange curves in the upper left panel. In contrast, for the 2:1 population, these errors appear to be more significant, causing some overlapping of the shaded areas (color-coded in the same way as for the 3:2 case) corresponding to the evolution of the SBDB population with and without massive Twotinos. We also present the cases, including Pluto alone with the 2:1 population, as well as Pluto together with the ten massive Twotinos (green curve, a case only applicable to the 2:1 populations). The lower panels show the same cases, color-coded in the same way as the upper panels, but for the L7 populations.
@@
Figure 7: Evolution of normalized counts of resonant objects from SBDB (upper panels) and L7 (lower panels) samples. The left panels correspond to simulations of the 3:2 population, and the right panels to the 2:1 population. The solid color lines indicate the number of objects, while the dotted lines correspond to exponential fits as described in equation 6 and table 3 in Section IV.4. Normalization is done with respect to the number of objects at 10 Myr. The colored shadows in the upper panels represent the variations obtained from the 3 different runs we performed in each scenario for the SBDB sample. Lines corresponding to the L7 model (bottom panels) do not have shadows, as their simulations were conducted only once due to computational expense.
In the case of Twotinos (right panels of Fig. 7), the evolutionary track of simulations that included only the giant planets is very similar for both the SBDB and L7 populations (blue curves in the upper and lower right panels). Adding the ten massive Twotinos into the simulations (orange curves) results in only a slight difference compared to the giant planet case alone, i.e., the perturbations produced by the most massive Twotinos are almost negligible, as were the perturbations of the nine massive Plutinos besides Pluto for the 3:2 population. Surprisingly, a more critical instability on the 2:1 population is caused by Pluto alone, as its presence has a bigger effect than the ten massive Twotinos. This can be observed by comparing the red and green curves with the orange curve, and this effect is also evident in the L7 model, as shown by the corresponding curves in the bottom-right panel.
Pluto’s influence on the 2:1 populations may be attributed to the dwarf planet’s close encounters at aphelion with Twotino bodies. However, this suggests that Eris might also play an important role for Plutino bodies, given its perihelion is approximately 38 au. To explore this idea, we conducted numerical simulations using the 3:2 populations from the SBDB and L7 samples under perturbations from the giant planets and Eris. Our findings indicate that Eris has a minimal impact on the leak rate of the resonance, accounting for approximately 1% more leaking, when compared to the case of the giant planets alone. Another possibility for Pluto’s effect on the 2:1 population is their mutual 4:3 MMR. The significance of this resonance has not been previously explored, since Pluto is not massive enough to trap objects in its resonances. However, Neptune can trap Twotinos in Pluto’s 4:3 resonance, where Pluto can perturb them secularly. We are exploring this idea in Ramírez-Vargas et al. in prep.
IV.4 Leaking rate fittings and comparison with previous works
@@
where $N_{T}(t)$ is the resonant population at time $t$, which can be divided into two populations, $N_{S}$ and $N_{U}$ representing the stable and unstable populations, respectively, with $N_{U}^{0}$ being the unstable population at $t=0$. Note that $N_{S}$ is a constant offset related to the stable fraction of particles in the resonance, and $\tau$ gives the exponential decay rate of the unstable population. This behavior is consistent with the stochastic nature of the leaking phenomenon from resonances, i.e., the escape of objects from the resonances appears to be driven by random processes, as indicated by this distribution. At the same time, the constant offset can be attributed to hyper-stable objects, such as Pluto (e.g. Ito and Malhotra, 2025), but it consists of a significant fraction of the original population.
@@
Though not many, there have been some previous quantifications of the leaking rates from Neptune’s 3:2 and 2:1 MMRs. Morbidelli (1997) estimates a power law function of the time with an exponent of -0.5 for the 3:2 MMR, while Tiscareno and Malhotra (2009) found exponents of -0.55 and -0.77 for the 3:2 and 2:1 MMRs, respectively. Based on these works, Greenstreet et al. (2015) estimated the impact rates on Pluto and Charon during the last 4 Gyr, using a power law fitting given by:
@@
with $b=0.52$ for the 3:2 MMR and $b=0.77$ for the 2:1 MMR. Greenstreet et al. (2015) defined equation 8 so that the present time corresponds to 4.5 Gyr; therefore, to use it, we have to extrapolate it from 4.5 to 8.5 Gyr (i.e., we use it from the present time to 4 Gyr into the future).
In Fig. 8 we compare Eq. 8 with our exponential fittings for each resonance and for both the SBDB and L7 populations. Although Greenstreet et al. (2015) models only consider the four giant planets as perturbers, in this comparison, we include simulations with only the giant planets as perturbers, as well as our most complete simulations; that is, in the case of Plutinos, the case including the ten massive Plutinos, and for Twotinos, the case including the ten massive Twotinos and Pluto.
Overall, we only find agreement with our SBDB simulations without massive Plutinos and Twotinos (blue lines in both panels), having a very close match for the 2:1 MMR, and a good match to the initial slope in the 3:2 resonance. Since the real Trans-Neptunian region contains massive objects, it stands to reason that the erosion of the MMRs is faster than what our GP-only simulations or Greenstreet equations imply, suggesting that the MMRs were more heavily populated than the Greenstreet model predicts, at least for the past 1 Gyr (based on a short extrapolation of our equations into the past). This in turn has implications for the age of Pluto’s (and Charon’s) surface, with our model suggesting younger surfaces than predicted by Greenstreet et al. (2015).
Figure 8: Leaking rate fit comparisons. Our rates (solid colored lines) come from Table 3, but have been renormalized so they start with a value of 1.00 at $t=0$ Gyr. The dashed line corresponds to Eq. 8, taken from Greenstreet et al. (2015).
IV.5 Effect of massive Plutinos and Twotinos on the evolution of resonant populations
@@ V Conclusions
In section IV.1, we conducted an analysis of the final state of the simulations that include only the GPs after 4 Gyr of integration. The sculpting of the orbital parameters tends to converge for the SBDB and the L7 populations, especially in semimajor axis. However, the median inclination of the 3:2 population, as well as the median eccentricity of the 2:1 population, did not really converge, indicating that no amount of evolution will make these distributions consistent among each other. The SBDB sample has observational biases, and whether there are inclination over or underestimations in the L7 model is beyond the scope of this paper. We also note that larger inclinations facilitate the escape from resonance, thus particles in the large-$i$ tail leak faster; this results in the median inclinations of all the populations being smaller at 4 Gyr.
The comparison between the initial and final distribution of the libration amplitudes and libration centers revealed that asymmetries in the leading and trailing populations of the 2:1 population remain unchanged over the evolution of our simulations: L7 objects, which start with a symmetrical distribution, end up with the same rough amount of leading and trailing objects, and the SBDB objects, which start with an uneven distribution among libration centers, ends up with a different, still uneven distribution at the 4 Gyr mark. This suggests that asymmetries in the leading and trailing populations probably arise due to a secondary mechanism at the early stage of the formation of the original populations, most commonly attributed to Neptune’s migration at the early stages of the Solar System (e.g., Nesvorný, 2018, and references therein).
By adding the massive perturbers to the simulations, as discussed in Section IV.2, we explored the effect of these additional perturbations among themselves, as well as over the resonant populations as a whole. Regarding the massive objects, from the ten most massive Plutinos, we found that when only the GPs are considered, all ten objects are resonant for the whole 4 Gyr, while adding mass results in 9 being resonant for 4 Gyr, while the remaining one, $2003\,{\rm UZ}_{413}$, is resonant for about 1.5 Gyr. For the massive Twotinos, even when only the GPs are included as perturbers, only two massive Twotinos are resonant for the whole 4 Gyr (although a third one is recaptured near the end of the simulation); the other 8 leave the resonance at different times between 20 Myr and 3.2 Gyr. When adding mass to these Twotinos, they become more stable, with six remaining resonant for the 4 Gyr (and one more being recaptured). Yet when adding Pluto as a massive TNO, the set becomes less stable with only four massive Twotinos remaining resonant for the 4 Gyr (and, once more, one more being recaptured).
@@
The addition of massive bodies into the simulations had a prominent effect on each group, drastically diminishing their stability in some cases, particularly for the 3:2 population. Within the 4 Gyr of our simulations, Pluto is responsible for more than 24% of the objects escaping the 3:2 population, while only 2% is attributed to the other nine most massive Plutinos. The ten most massive Twotinos create a difference of 4% in the number of objects leaving the resonance. The most surprising feature found is the influence of Pluto on the 2:1 population, where 13% of the escapes are caused by the addition of Pluto alone. The influence of Pluto over the stability of the 2:1 MMR is quite unexpected. It could be attributed to perturbations from the dwarf planet when it lies closer to aphelion, when its distance to objects trapped in 2:1 MMR could be drastically reduced, increasing the interaction probability, but if such were the case, Eris should be as relevant for the 3:2 population stability given its perihelion of $\sim 38.4$ au. Preliminary test simulations showed that Eris does not significantly perturb the 3:2 population; thus, a different possibility is that Pluto affects the 2:1 MMR populations through their mutual 4:3 MMR. A more in-depth study of this phenomenon is left for future work (Ramírez-Vargas et al. in prep.).
When comparing our results with those of previous authors, we found that simulations of the SBDB without massive perturbers are mostly consistent with extrapolations of the theoretical models (e.g. Greenstreet et al., 2015), however, our simulations including massive TNOs result in greater leaking rates.
@@
We acknowledge the referee, B. Gladman, for a careful reading and report that helped improve this paper. This research was performed using services/resources provided by Grid UNAM, which is a collaborative effort driven by DGTIC and the research institutes of Astronomy, Nuclear Science and Atmosphere Science and Climate Change at UNAM. M.A.M. acknowledges Universidad de Atacama for the DIUDA grant No. 88231R14. A.P.-V. acknowledges the DGAPA-PAPIIT grants IA103224 and IN112526.
{contribution}
MAMG and AP came up with the initial research concept, most of the writing, and most of the editing. SR was responsible for the computer simulations, as well as some writing and editing. APV and CP contributed to reviewing, writing, and editing.